# SITE: towards Spatial Intelligence Thorough Evaluation

### v1

@@ 4 Data
We adopt a two-stage approach to constructing SITE as shown in Figure 2. To begin, in Section 4.1, we compile a list of evaluation tasks from a suite of benchmarks that are focused on evaluating existing AI model’s SI. Through a series of analysis and filtering steps, we construct a unified multiple-choice QA benchmark in a bid to standardize evaluations of SI in VLMs. Next, using the empirical analysis performed earlier as well as the the classification framework introduced in Section 3, we identify two underrepresented aspects of SI that are largely unaddressed by existing spatial reasoning benchmarks. To address these gaps, we propose new evaluation tasks to enable a more comprehensive and holistic evaluation of VLMs’ spatial reasoning capabilities in Section 4.2. Finally, we describe the curation of our final SITE dataset in Section 4.3.
@@ 4.1 Data Collection
Given the large number of isolated and fragmented evaluation benchmarks that focus on different aspects of SI, collecting a large-scale and comprehensive evaluation benchmark presents a significant challenge. With this challenge in mind, we propose a systematic data collection pipeline designed to filter and sample tasks relevant to SI in a structured manner as Figure 2.
@@
Data collection and filtering. To begin, we manually select a set of representative VQA datasets that potentially cover some aspects of SI (Figure 2). The initial collection of datasets comprise 22 image datasets and 8 video datasets, where examples of the former include VSR [43] as well as CV-Bench [64] and the latter include VSI-Bench [74] and VideoMME [18], respectively. We provide a detailed description of our selected benchmarks in the supplemental material. We note that we focus exclusively on the validation and test splits for each dataset. While the aforementioned datasets are generally large, they often only contain a subset of evaluation samples that are related to SI. Thus, it is necessary for us to filter out evaluation samples that are not relevant to the goal of evaluating spatial reasoning in VLMs. To filter out irrelevant evaluation samples, we adopt different strategies based on the existing annotations in each dataset. For datasets that contain category labels for the evaluations samples, we only retain evaluation samples that fall under the categories are relevant to spatial reasoning while discarding the rest. For datasets without predefined labels, we use a pretrained Large Language Model (LLM) to perform filtering of their constituent evaluation samples as described below.
@@
LLM filtering. To further refine the quality and relevancy of the collected evaluation samples, we employ a two-stage filtering process by leveraging GPT-4o [55], a powerful LLM, to classify the evaluation samples (Figure 2). A major issue we faced in curating a comprehensive spatial reasoning benchmark is that different datasets use different labels for their evaluation samples, even though they might be evaluating about similar aspects of SI. We tackle this issue by prompting the LLM to generate 6 coarse categories of tasks pertaining to SI. Specifically, we amass a set of original question labels from existing datasets and create a prompt for the LLM along with one to two sample datapoints as context for generating the task categories. The six coarse-level spatial intelligence categories are as follows: Counting and Existence (Count.), Spatial Relationship Reasoning (Rel.), Object Localization and Positioning (Loc.), 3D Information Understanding (3D Inf.), Multi-View Reasoning (MultiV.), and Movement Prediction and Navigation (Mov.) We proceed by classifying the valid spatial intelligence evaluation samples across all datasets into the abovementioned six categories.
@@
Despite this categorization, many evaluation samples lack task type labels, or have noisy annotations. To refine the classification, we conduct a second filtering stage using GPT-4o. We design a prompt template by incorporating key textual dataset information from each evaluation sample (e.g., questions, answers, options, and descriptions) along with carefully selected example data. The LLM is then queried to determine: (1) Whether the task pertains to spatial intelligence. (2) If so, which coarse category it falls under. We provide an example of the prompt template in the supplemental material.
@@
Statistics and Reform QA types. After undergoing LLM-based filtering, the resulting dataset is reduced to 223,083 task examples, comprising 206,887 image-based and 16,196 video-based QA pairs relevant to spatial intelligence. We conduct a statistical analysis of the data across the coarse-level spatial categories and observe a significant data imbalance. Specifically, Relationship Reasoning (Rel.) and Counting and Existence (Count.) problems dominate the evaluation. In contrast, tasks such as Multi-View Reasoning (MultiV.) is underrepresented to a large degree. Additionally, due to the diverse data sources, the QA formats can vary considerably across different datasets. To ensure consistency and ease of evaluation, we standardize all tasks to a multiple-choice QA format, where we reformulate open-ended QA tasks accordingly.
@@ 4.2 Novel Proposed Tasks
To apply the cognitive science-based classification framework introduced in Section 3, we also conduct a fine-grained manual annotation of the filtered SI tasks. Following the classification workflow illustrated in Figure 2, we obtained the distribution of the collected dataset under different classification systems, as provided in the appendix III. Our analysis reveals that there is a significant lack of tasks involving perspective transformations, which indicates that most existing SI evaluation tasks are constrained to reasoning from a fixed camera viewpoint. However, perspective transformation is especially essential for spatial reasoning in real-world scenarios such as navigation and route planning, where humans are naturally able to interpret spatial relationships from multiple viewpoints. To bridge this gap, we introduce two novel tasks specifically designed to evaluate extrinsic-static and extrinsic-dynamic spatial reasoning under perspective-transformed conditions.
Figure 3: Ego-Exo view association tasks. The goal of this task is pick the correct exocentric view given the egocentric view of a visual scene or vice versa.
Data Preparation. To assess models’ perspective transformation abilities in real-world scenarios, we leverage videos and annotations from the EgoExo4D [22] dataset. EgoExo4D provides 5,035 takes across eight real-world scenarios, where each take includes at least one egocentric view and multiple exocentric views captured synchronously. This multi-view setting provides a strong foundation for our task design, enabling a robust evaluation of spatial intelligence across diverse viewpoints.
Shuffled Frames Reordering. In this extrinsic-dynamic task, as shown in Figure 8, a model has to infer the correct temporal order from multiple viewpoints. Specifically, given a video clip, we extract the start and end frames from the egocentric video as reference points. Within this segment, we also sample 4 frames that capture key motion events from the exocentric views before shuffling them randomly. Finally, the model must predict the correct sequence by reasoning about motion dynamics across space and time. This task is also evaluated in reverse, by requiring models to predict the correct temporal order of egocentric frames based on the provided exocentric views. To ensure interpretability and feasibility of our proposed frame-reordering task, we also incorporate another round of human annotations to remove ambiguous and extremely difficult cases.
4.3 SITE
To ensure a balanced representation of various SI factors, we perform stratified sampling on the collected data from Section 4.1 to achieve an even distribution across different spatial categories while maintaining diversity. Including the two newly proposed tasks, our final benchmark consists of 8,068 QA pairs, including 4,260 image-based QA pairs and 3,808 video-based QA pairs, covering 31 computer vision datasets. Table 2 shows some statistics.
5 Experiments
@@ 5.1 Benchmark Evaluation
Evaluation Models. We evaluate 9 state-of-the-art VLMs that accept both image and video inputs, covering a diverse range of model architectures and parameter scales. From open-sourced models, we select LLAVA-OneVision [39], InternVL-2.5 [9], Qwen2.5-VL [73], and Phi-3.5V [1]. For proprietary models, we evaluate GPT-4o [55] and Gemini 1.5 [62]. To ensure standardized and reproducible evaluation, we utilize the lmms-eval [82, 38] framework for benchmarking all models.
@@
where $N$, $n_{i}$ and $X_{i}$ are the total number of questions, number of answer choices for the $i$-th multiple-choice question and an indicator variable, respectively. If the model correctly answer the $i$-th question, we set $X_{i}=1$. Otherwise, we set $X_{i}=0$. $\mathcal{CAA}=1$ when all predictions are correct ($X_{i}=1$ for all $i$), indicating perfect performance; $\mathcal{CAA}=0$ when the preditions perform no better than random guessing ($\sum X_{i}=\sum_{i=1}^{N}\frac{1}{n_{i}}$); $\mathcal{CAA}<0$ reflects performace worse than random guess. This adjustment ensures that model performance is evaluated as the improvement beyond random chance, providing a fairer comparison across questions with different numbers of answer options.
@@ 5.2 Results on SITE
We begin by reporting the results of our evaluation on our proposed tasks of view association and frames reordering in Table 3. To benchmark the difficulty of these two tasks, we conduct a human study and the results demonstrate that humans are actually very adept at understanding and reasoning about 3D visual scenes from both egocentric and exocentric viewpoints. We randomly sampled 30 questions from each of the two views and found that human participants achieved perfect accuracy (100%). However, even state-of-the-art proprietary and open-sourced models such as GPT-4o and InternVL-2.5-8B achieve very low performances compared to the human level. This trend is also corroborated by the accuracy obtained by the abovementioned models on the full split of the view association task where GPT-4o and Gemini-1.5-pro achieves an average of 28.20% and 7.03%, respectively. Interestingly, while VLMs such as LLaVA-OneVision and Qwen2.5-VL have been demonstrated to perform well on conventional image and video question answering benchmarks like MME and VideoMME, their perception capabilities do not translate well to reasoning about spatial relationships and information beyond just perception. For instance, we observe that highly capable VLMs such as LLAVA-OneVision-7B and QWen2.5-VL-7B achieve very low average $\mathcal{CAA}$ of 15.95% and 0.65% on this view association task. We hypothesize that the lack of training data involving viewpoint transformations in large vision-language datasets—where most reasoning tasks rely directly on the camera’s perspective—is the primary cause. The severe data imbalance observed during our dataset collection further supports this hypothesis.
@@
Frames Reordering. Furthermore, we also observe similar performance trend on our proposed task of temporal frames reordering (Table 3 right). Once again, the results show that humans are able to understand the temporal occurrence of events from different viewpoints, where they achieve close to perfect accuracy of 98% and 96% in the ego2exo and exo2ego directions, respectively. Interestingly, the GPT-4o model experiences a sharp drop in performance on this task, as compared to the task of view association. This might suggest that the GPT-4o model is not able to understand the mapping between different viewpoints of temporal events. Additionally, it is notable that the large-scale and proprietary models are severely underperforming open-sourced VLMs such as InternVL-2.5-8B and QWen2.5-VL-7B. In fact, the Qwen2.5-VL-7B model achieves the best average performance of 5.93%. One possible reason underlying this result is that the Qwen2.5-VL model is trained on video grounding data, which helps the model to learn a more effective understanding of time, and consequently temporal order of events, in videos.
Why do humans excel while VLMs fail? The tasks of view association and frames ordering explicitly evaluate perspective transformation and temporal reasoning in spatial intelligence. Humans might excel at them since we can effortlessly imagine a visual scene from different viewpoints, leveraging lived experiences and real-world knowledge to establish correspondences between different viewpoints. In contrast, VLMs may fail to understand the spatial mapping between egocentric and exocentric views of the same visual scenes, relying on pixel-level similarities rather than geometric understanding and temporal dynamics.
@@
Evaluation on the full SITE. We report our evaluation of state-of-the-art open-sourced and proprietary models in Table 4. For more detailed analysis, we break down the results of the evaluation across the six coarse categories of spatial intelligence tasks, as discussed in Section 3. To begin, the performance achieved by various open-sourced and proprietary models are consistent with our observations in Table 3. There is a large performance gap between the accuracy obtained by humans and state-of-the-art VLMs, which suggests that simply scaling up the amount of supervised finetuning (SFT) and instruction following multimodal data for pretraining may be inadequate in helping these VLMs to acquire effective spatial intelligence. It is also notable that humans only achieve an overall $\mathcal{CAA}$ scores of 67.5%, which hints at the difficulty of our proposed SITE. Interestingly, humans perform significantly worse on counting(e.g., counting objects in a long video), 3D understanding(e.g., inferring camera’s transformation matrix), and movement prediction(e.g., navigating within a scene presented in a long video) as compared to other three categories. This result might be due to humans’ attention bottlenecks in tracking multiple objects [16], explaining why counting moving objects and tracking spatial transformations across time is very challenging.
@@
However, the best-performing VLM GPT-4o still underperforms the human performance by $\sim$32%. The Qwen2.5-VL-7B and InternVL-2.5-8B models lead in overall accuracy with 31.4% and 32.8%, respectively. Despite containing much fewer model parameters, these open-sourced models perform competitively with GPT-4o and Gemini-1.5-Pro. Interestingly, we observe that Qwen2.5-VL-7B performs the best among all open-sourced VLMs on localization. This might be due to the pretraining recipe of Qwen2.5-VL-7B which also includes image and video grounding tasks [73]. We also see that multi-view reasoning is especially challenging for VLMs in general, where the performance obtained by GPT-4o is lower than that of human performance by over 70% on the tiny subset. On the full split, all of the state-of-the-art VLMs obtain $\mathcal{CAA}$ of less than 10%. One possible reason for the low performance is that these VLMs are generally not trained with different viewpoints for the same image or video.
@@
From these empirical results, we also observe the benefits of using models with larger numbers of parameters. As evidenced by the consistent performance gains obtained by LLaVA-OneVision-7B and Qwen2.5-VL-7B over their smaller counterparts, VLMs with a higher number of parameters generally are able to perform spatial reasoning of visual scenes and environments much more effectively. However, it is notable that 3D understanding scores are consistently low across all VLMs, with most models scoring below 15%, indicating a persistent challenge in understanding depth(e.g., reasoning the depth relationships between objects) and three-dimensional spatial transformations(e.g., folding a 2D grid into a cube).

### v2

@@ 4 Data
We adopt a two-stage approach to constructing SITE. To begin, in Section 4.1, we compile a list of evaluation tasks from a suite of benchmarks that are focused on evaluating existing AI models’ SI. Through a series of analysis and filtering steps, we construct a unified multiple-choice QA benchmark in a bid to standardize evaluations of SI in VLMs. Next, using the empirical analysis performed earlier as well as the classification framework introduced in Section 3, we identify two underrepresented aspects of SI that are largely unaddressed by existing spatial reasoning benchmarks. To address these gaps, we propose new evaluation tasks to enable a more comprehensive and holistic evaluation of VLMs’ spatial reasoning capabilities in Section 4.2. Finally, we describe the curation of our final SITE dataset in Section 4.3.
@@ 4.1 Data Collection
Given the large number of isolated and fragmented evaluation benchmarks that focus on different aspects of SI, collecting a large-scale and comprehensive evaluation benchmark presents a significant challenge. With this challenge in mind, we propose a systematic data collection pipeline designed to filter and sample tasks relevant to SI in a structured manner, as Figure 2.
@@
Data collection and filtering. To begin, we manually select a set of representative VQA datasets that potentially cover some aspects of SI (Figure 2). The initial collection of datasets comprises 22 image datasets and 8 video datasets, where examples of the former include VSR [Liu2022VisualSR] as well as CV-Bench [tong2024cambrian1], and the latter include VSI-Bench [yang2024thinkingspacemultimodallarge] and VideoMME [fu2024video], respectively. We provide a detailed description of our selected benchmarks in the supplemental material. We note that we focus exclusively on the validation and test splits for each dataset. While the aforementioned datasets are generally large, they often only contain a subset of evaluation samples that are related to SI. Thus, it is necessary for us to filter out evaluation samples that are not relevant to the goal of evaluating spatial reasoning in VLMs. To filter out irrelevant evaluation samples, we adopt different strategies based on the existing annotations in each dataset. For datasets that contain category labels for the evaluation samples, we only retain evaluation samples that fall under categories relevant to spatial reasoning while discarding the rest. For datasets without predefined labels, we use a pretrained Large Language Model (LLM) to perform filtering of their constituent evaluation samples as described below.
@@
LLM filtering. To further refine the quality and relevancy of the collected evaluation samples, we employ a filtering process by leveraging GPT-4o [openai2024gpt4technicalreport], a powerful LLM, to classify the evaluation samples (Figure 2). A major issue we faced in curating a comprehensive spatial reasoning benchmark is that different datasets use different labels for their evaluation samples, even though they might be evaluating similar aspects of SI. We tackle this issue by prompting the LLM to generate 6 coarse categories of tasks pertaining to SI. Specifically, we amass a set of original question labels from existing datasets and create a prompt for the LLM along with one to two sample datapoints as context for generating the task categories. The six coarse-level spatial intelligence categories are as follows: Counting and Existence (Count.), Spatial Relationship Reasoning (Rel.), Object Localization and Positioning (Loc.), 3D Information Understanding (3D Inf.), Multi-View Reasoning (MultiV.), and Movement Prediction and Navigation (Mov.). We proceed by classifying the valid spatial intelligence evaluation samples across all datasets into the abovementioned six categories.
@@
Despite this categorization, many evaluation samples lack task-type labels or have noisy annotations. To refine the classification, we conduct a filtering stage using GPT-4o. We design a prompt template by incorporating key textual dataset information from each evaluation sample (e.g., questions, answers, options, and descriptions) along with carefully selected example data. The LLM is then queried to determine: (1) Whether the task pertains to spatial intelligence; (2) If so, which coarse category it falls under. We provide an example of the prompt template in the supplemental material.
@@
Statistics and Reform QA types. After undergoing LLM-based filtering, the resulting dataset is reduced to 223,083 task examples, comprising 206,887 image-based and 16,196 video-based QA pairs relevant to spatial intelligence. We conduct a statistical analysis of the data across the coarse-level spatial categories and observe a significant data imbalance. Specifically, Relationship Reasoning (Rel.) and Counting and Existence (Count.) problems dominate the evaluation. In contrast, tasks such as Multi-View Reasoning (MultiV.) are underrepresented to a large degree. Additionally, due to the diverse data sources, the QA formats can vary considerably across different datasets. To ensure consistency and ease of evaluation, we standardize all tasks to a multiple-choice QA format, where we reformulate open-ended QA tasks accordingly.
@@ 4.2 Novel Proposed Tasks
To apply the cognitive science-based classification framework introduced in Section 3, we also conduct a fine-grained manual annotation of the filtered SI tasks. Following the classification workflow illustrated in Figure 2, we obtained the distribution of the collected dataset under different classification systems, as provided in the appendix. Our analysis reveals that there is a significant lack of tasks involving perspective transformations, which indicates that most existing SI evaluation tasks are constrained to reasoning from a fixed camera viewpoint. However, perspective transformation is especially essential for spatial reasoning in real-world scenarios such as navigation and route planning, where humans are naturally able to interpret spatial relationships from multiple viewpoints. To bridge this gap, we introduce two novel tasks specifically designed to evaluate extrinsic-static and extrinsic-dynamic spatial reasoning under perspective-transformed conditions.
Figure 3: Ego-Exo view association tasks. The goal of this task is to pick the correct exocentric view given the egocentric view of a visual scene or vice versa.
Data Preparation. To assess models’ perspective transformation abilities in real-world scenarios, we leverage videos and annotations from the Ego-Exo4D [grauman2024egoexo4dunderstandingskilledhuman] dataset. Ego-Exo4D provides 5,035 takes across eight real-world scenarios, where each take includes at least one egocentric view and multiple exocentric views captured synchronously. This multi-view setting provides a strong foundation for our task design, enabling a robust evaluation of spatial intelligence across diverse viewpoints.
@@
Shuffled Frames Reordering. In this extrinsic-dynamic task, a model has to infer the correct temporal order from multiple viewpoints. Specifically, given a video clip, we extract the start and end frames from the egocentric video as reference points. Within this segment, we also sample 4 frames that capture key motion events from the exocentric views before shuffling them randomly. Finally, the model must predict the correct sequence by reasoning about motion dynamics across space and time. This task is also evaluated in reverse, by requiring models to predict the correct temporal order of egocentric frames based on the provided exocentric views. To ensure interpretability and feasibility of our proposed frame-reordering task, we also incorporate another round of human annotations to remove ambiguous and extremely difficult cases.
4.3 SITE
To ensure a balanced representation of various SI factors, we perform stratified sampling on the collected data from Section 4.1 to achieve an even distribution across different spatial categories while maintaining diversity. Including the two newly proposed tasks, our final benchmark consists of 8,068 QA pairs, including 4,260 image-based QA pairs and 3,808 video-based QA pairs, covering 30 existing benchmark datasets and the Ego-Exo4D dataset. Table 2 shows some statistics.
5 Experiments
@@ 5.1 Benchmark Evaluation
Evaluation Models. We evaluate 9 state-of-the-art VLMs that accept both image and video inputs, covering a diverse range of model architectures and parameter scales. From open-sourced models, we select LLAVA-OneVision [li2024llavaonevisioneasyvisualtask], InternVL-2.5 [chen2024internvl], Qwen2.5-VL [qwen2.5-VL], and Phi-3.5V [abdin2024phi3technicalreporthighly]. For proprietary models, we evaluate GPT-4o [openai2024gpt4technicalreport] and Gemini 1.5 [gemini_1_5]. To ensure standardized and reproducible evaluation, we utilize the lmms-eval [lmms_eval2024] framework for benchmarking all models.
@@
where $N$, $n_{i}$, and $X_{i}$ are the total number of questions, number of answer choices for the $i$-th multiple-choice question, and an indicator variable, respectively. If the model correctly answers the $i$-th question, we set $X_{i}=1$. Otherwise, we set $X_{i}=0$. $\mathcal{CAA}=1$ when all predictions are correct ($X_{i}=1$ for all $i$), indicating perfect performance; $\mathcal{CAA}=0$ when the preditions perform no better than random guessing ($\sum X_{i}=\sum_{i=1}^{N}\frac{1}{n_{i}}$); $\mathcal{CAA}<0$ reflects performace worse than random guess. This adjustment ensures that model performance is evaluated as the improvement beyond random chance, providing a fairer comparison across questions with different numbers of answer options.
@@ 5.2 Results on SITE
We begin by reporting the results of our evaluation on our proposed tasks of view association and frames reordering in Table 3. To benchmark the difficulty of these two tasks, we conduct a human study and the results demonstrate that humans are actually very adept at understanding and reasoning about 3D visual scenes from both egocentric and exocentric viewpoints. We randomly sampled 30 questions from each of the two views and found that human participants achieved perfect accuracy (100%). However, even state-of-the-art proprietary and open-sourced models such as GPT-4o and InternVL-2.5-8B achieve very low performances compared to the human level. This trend is also corroborated by the accuracy obtained by the abovementioned models on the full split of the view association task where GPT-4o and Gemini-1.5-pro achieve an average of 28.20% and 7.03%, respectively. Interestingly, while VLMs such as LLaVA-OneVision and Qwen2.5-VL have been demonstrated to perform well on conventional image and video question answering benchmarks like MME and VideoMME, their perception capabilities do not translate well to reasoning about spatial relationships and information beyond just perception. For instance, we observe that highly capable VLMs such as LLAVA-OneVision-7B and QWen2.5-VL-7B achieve very low average $\mathcal{CAA}$ of 15.95% and 0.65% on this view association task. We hypothesize that the lack of training data involving viewpoint transformations in large vision-language datasets—where most reasoning tasks rely directly on the camera’s perspective—is the primary cause. The severe data imbalance observed during our dataset collection further supports this hypothesis.
@@
Frames Reordering. Furthermore, we also observe a similar performance trend on our proposed task of temporal frames reordering (Table 3 right). Once again, the results show that humans are able to understand the temporal occurrence of events from different viewpoints, where they achieve close to perfect accuracy of 98% and 96% in the ego2exo and exo2ego directions, respectively. Interestingly, the GPT-4o model experiences a sharp drop in performance on this task, as compared to the task of view association. This might suggest that the GPT-4o model is not able to understand the mapping between different viewpoints of temporal events. Additionally, it is notable that the large-scale and proprietary models are severely underperforming open-sourced VLMs such as InternVL-2.5-8B and QWen2.5-VL-7B. In fact, the Qwen2.5-VL-7B model achieves the best average performance of 5.93%. One possible reason underlying this result is that the Qwen2.5-VL model is trained on video grounding data, which helps the model to learn a more effective understanding of time and consequently, the temporal order of events in videos.
@@
Evaluation on the full SITE. We report our evaluation of state-of-the-art open-sourced and proprietary models in Table 4. For more detailed analysis, we break down the results of the evaluation across the six coarse categories of spatial intelligence tasks, as discussed in Section 3. To begin, the performance achieved by various open-sourced and proprietary models is consistent with our observations in Table 3. There is a large performance gap between the accuracy obtained by humans and state-of-the-art VLMs, which suggests that simply scaling up the amount of supervised fine-tuning (SFT) and instruction following multimodal data for pretraining may be inadequate in helping these VLMs to acquire effective spatial intelligence. It is also notable that humans only achieve an overall $\mathcal{CAA}$ score of 67.5%, which hints at the difficulty of our proposed SITE. Interestingly, humans perform significantly worse on counting(e.g., counting in a long video), 3D understanding(e.g., inferring the camera’s transformation matrix), and movement prediction(e.g., navigating in a long video) as compared to the other three categories. This result might be due to humans’ attention bottlenecks in tracking multiple objects [franconeri2013flexible], explaining why counting moving objects and tracking spatial transformations across time is very challenging.
@@
However, the best-performing VLM GPT-4o still underperforms the human performance by $\sim$32%. The Qwen2.5-VL-7B and InternVL-2.5-8B models lead in overall accuracy with 31.4% and 32.8%, respectively. Despite containing much fewer model parameters, these open-sourced models perform competitively with GPT-4o and Gemini-1.5-Pro. Interestingly, we observe that Qwen2.5-VL-7B performs the best among all open-sourced VLMs on localization. This might be due to the pretraining recipe of Qwen2.5-VL-7B which also includes image and video grounding tasks [qwen2.5-VL]. We also see that multi-view reasoning is especially challenging for VLMs in general, where the performance obtained by GPT-4o is lower than that of human performance by over 70% on the tiny subset. On the full split, all of the state-of-the-art VLMs obtain $\mathcal{CAA}$ of less than 10%. One possible reason for the low performance is that these VLMs are generally not trained with different viewpoints for the same image or video.
@@
From these empirical results, we also observe the benefits of using larger models. As evidenced by the consistent performance gains obtained by LLaVA-OneVision-7B and Qwen2.5-VL-7B over their smaller counterparts, VLMs with a higher number of parameters generally are able to perform spatial reasoning of visual scenes and environments much more effectively. However, it is notable that 3D understanding scores are consistently low across all VLMs, with most models scoring below 15%, indicating a persistent challenge in understanding depth(e.g., reasoning the depth relationships between objects) and three-dimensional spatial transformations(e.g., folding a 2D grid into a cube).